OpenStudy (anonymous):

If f(x)=x^2-7 and h(x)=7x-5, then f(h(2))

OpenStudy (anonymous):

what is \(h(2)\)? you need that first

OpenStudy (anonymous):

so 14x-25?

OpenStudy (unklerhaukus):

h(x) = 7x-5 h(2) = 7(2)-5 = (should not have x's in it)

OpenStudy (anonymous):

oh no lets go slow

OpenStudy (anonymous):

okay so I have 9 now

OpenStudy (unklerhaukus):

that's better h(2) =9 now what is f(h(2))=f(9) remember f(x)=x^2-7 so sub in x=9

OpenStudy (anonymous):

74?

OpenStudy (unklerhaukus):

yep